i've never encountered a question about solving the system that involved X^2 and y^2. I'm confused on how to first tackle the problem.
3y-x=5
x^2-y^2=-3

Answer by Alan3354(69443)

3y-x=5 --> x = 3y - 5
x^2-y^2=-3
-------------
Sub for x in the 2nd eqn
(3y-5)^2 - y^2 = -3
9y^2 - 30y + 25 - y^2 = -3
8y^2 - 30y + 28 = 0
4y^2 - 15y + 14 = 0
(y - 2)*(4y - 7) = 0
y = 2
y = 7/4
--------
Sub and solve for x.